Write the standard form of the equation of the line through the given point with the given slope.
1. Through: (-5,0), slope = -3/5

Answer by stanbon(75887)

Write the standard form of the equation of the line through the given point with the given slope.
Through: (-5,0), slope = -3/5
-------------------
y = mx + b
0 = (-3/5)(-5) + b
--------
0 = 3 + b
b = -3
-------
Equation:
y = (-3/5)x - 3
-----------------
Cheers,
Stan H.
------------------